Circular No. 8677

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006ak AND 2006al Two apparent supernovae have been discovered on unfiltered CCD

images: 2006ak by N. Tyurina, V. Lipunov, V. Kornilov, A. Krylov, G. Borisov, A. Belinski, D. Kuvshinov, E. Gorbovskoy, P. Gritsyk, and M. Sinitsyn (MASTER robotic telescope; cf. IAUC 8603); and 2006al by R. Holmes (Charleston, IL) and H. Devore (Fayetteville, NC) with a 0.40-m Schmidt-Cassegrain reflector.

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ak Feb. 17.97 11^{h}09^{m}32.^{s}83 +28^{o}37′50′′.3 15.8 6′′ W, 18′′ N
2006al Feb. 19.32 10 39 28.26 + 5 11 00.3 16.9 3′′.1 W, 0′′.7 S

SN 2006ak is a type-Ia supernova about a week past maximum light (cf. CBET 408). Additional approximate magnitudes of 2006ak: Feb. 8.99 UT, ∼ 17.0; 18.00, 15.8. Additional approximate magnitudes of 2006al: Jan. 26.29, [20.5; Feb. 21.24, 16.9 ± 0.3.

SUPERNOVAE 2006X AND 2006ab Unfiltered CCD survey magnitudes by Lipunov et al. for SN 2006X (cf.

IAUC 8667): Feb. 6.06 UT, 15.8; 8.96, 14.6.

D. S. Wong, J. M. Silverman, D. Pooley, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1020 nm), obtained on Feb. 22 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2006ab (IAUC 8669) is of type Ic, ∼ 2 weeks past maximum brightness. The Ca II near-infrared triplet absorption is especially strong.

COMETS 174P/ECHECLUS AND 175P/HERGENROTHER The Committee on Small Bodies Nomenclature has agreed to give the

comet P/2000 EC_{98} (cf. IAUC 8656, 8660) the same name as the centaur minor planet (60558), Echeclus (cf. MPC 55988), which has been assigned also the permanent comet number 174P (MPC 55911). P/2000 C1 = P/2006 A3 (Hergenrother) has been assigned the number 175P (cf. IAUC 8664).

RS OPHIUCHI Corrigenda. M. Bode advises us that the observation times given on

IAUC 8675 (lines 9 and 14 of the second item, printed version) should read Feb. 16.00 (or 3.17 days after the reported optical outburst) and 17.86 UT.

2006 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green